OpenStudy (anonymous):

really need help on geo2

OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

In circle O, CD = 44, OM = 20, ON = 19, , (The figure is not drawn to scale.) a. Find the radius. If your answer is not an integer, express it in radical form. b. Find FN. If your answer is not an integer, express it in radical form. c. Find EF. Express it as a decimal rounded to the nearest tenth.

OpenStudy (anonymous):

Where's the figure?

OpenStudy (anonymous):

i dont know how to add it

OpenStudy (anonymous):

Click on either Draw or Attach File below textbox.

OpenStudy (anonymous):

idk if i did this right

OpenStudy (anonymous):



OpenStudy (anonymous):

i drew it the best i could

OpenStudy (phi):

do we know anything else ?

OpenStudy (anonymous):

just what i posted i think i might know the answer to a but idk about b and c

OpenStudy (phi):

how do you solve for the radius ?

OpenStudy (anonymous):

a. 20^2 + 22^2 = x^2 400 + 484 = x^2 x^2 = 884 x = ??? x = sqrt(884) x = 2*sqrt(221)

OpenStudy (phi):

are you saying OM is perpendicular to CD ?

OpenStudy (anonymous):

thats my answer for a where they want the radius

OpenStudy (phi):